OpenStudy (mony01):

find the derivative f(x)=x^2lnx

OpenStudy (anonymous):



OpenStudy (australopithecus):

First split it into two functions

OpenStudy (australopithecus):

Future Math Prof if he wanted the answer he could just go to wolframalpha

OpenStudy (australopithecus):

so you have s(x) = x^2 and g(x) = ln(x) Derive both of these first then put it into the summation rule f'(x) = s(x)g'(x) + s'(x)g(x)

OpenStudy (anonymous):